A plate carries a charge of -3.9 µC, while a rod carries a charge of 2.5 µC. How many electrons must be transferred from the plate to the rod, so that both objects have the same charge?

N=?
Answer #1

Solution)

Here,

Q1 = -3.9 µC, Q2 = +2.5 µC

Let q = transferred negative charge

Q1 - q = Q2 + q

so q = (Q1 - Q2)/2 = -3.2 µC

Here,

number of electrons N = q/e, where e = -1.60*10^-19 C = -1.60*10^-13 µC

N = 3.2/1.60*10^-13 =2*10^13 (Ans)

=======

Good luck!:)